If 5 technicians can produce 64 surfboards in 4 days, how many surfboards can 8 technicians produce in 10 days?
Answer by stanbon(75887)

If 5 technicians can produce 64 surfboards in 4 days, how many surfboards can 8 technicians produce in 10 days?
--------
# of surfboards is directly related to # of days an # of technicians::
-----
s = k*d*t
-----
Solve for: k using "5 technicians can produce 64 surfboards in 4 days"
=============
64 = k*5*4
k = 64/20 = 3.2
--------
Equation:
s = 3.2*d*t
----------------
how many surfboards can 8 technicians produce in 10 days?
s = 3.2*10*8
-====
# of boards = 256
====================
Cheers,
Stan H.
===============